OpenStudy (anonymous):

PLEASE HELP MEEEEEEEEEEEEEEEEEEEEEEEEEE Elimination method. 5x + 2y = 55 2x - y = 19 answer is x+y=12 ...... I need to know why

OpenStudy (cj49):

are u sure x+y=12 is the ans

OpenStudy (cj49):

i got it

OpenStudy (cj49):

here if ulll substitute the values of x and y ull get 10 +2 =12 thats y the ans is x+y=12

OpenStudy (helder_edwin):

the answer is not x+y=12!!

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

ask me why

OpenStudy (cj49):

lol if u substitute the values it cms down to 10 +2=12(approximately)

OpenStudy (anonymous):

nope thats not why